### v7

# disposition

Abstract
Abstract. This paper develops a formal framework for analyzing information sufficiency in cooperative decision-making. Departing from models of incomplete information that emphasize accuracy of beliefs, we ask what kind and amount of information are required for justifying a choice. Using sequent calculus (Gentzen, 1935a, b), we construct a syntactic model that makes explicit an agent’s reasoning process: from which information, through what inference procedure, and to what conclusion. We define a syntactic criterion, called Ci-acceptability, to determine when a proposed distribution is justifiable given limited structural information. We show that, under a mild assumption that each coalition’s potential is known to at least one member, the set of unanimously accepted payoffs coincides with the core. This reinterprets the core not as a predictive solution, but as the boundary of what can be justified under minimal information. Moreover, our results offer a new perspective on Debreu and Scarf’s (1963) theorem: while the core converges to the competitive equilibrium in replicated markets, this convergence obscures a key asymmetry. Competitive equilibrium requires only local information—preferences and prices—regardless of market size, whereas reaching the core entails increasing informational demands that may ultimately exceed any agent’s cognitive capacity.
keywords:
Information sufficiency, Cooperative game theory, The core, Debreu-Scarf theorem, Proof theory
JEL Classification: C71, D51, D83, D90
Someone saw Nasrudin searching for something on the ground. “What have you lost, Mulla?” he asked. “My key,” said the Mulla. So they both went down on their knees and looked for it. After a time the other man asked: “Where exactly did you drop it?” “In my own house.” “Then why are you looking here?” “There is more light here than inside my own house.”—– “The exploits of the Incomparable Mulla Nasrudin” by Idries Shah, quoted in Doniger (2009)
1 Introduction
What kind of information—and how much of it—is sufficient for decision-making, particularly for justifying a proposed payoff distribution in a cooperative setting? This fundamental question not only lies at the heart of understanding decision-making under complicated circumstance where relevant information may be missed, it also has profound implications for how we interpret and apply canonical results in economic theory. Consider the classical Debreu-Scarf theorem (Debreu and Scarf, 1963), which demonstrates that the core converges to competitive equilibrium in replicated markets. From the aspect of information sufficiency, however, it obscures a crucial asymmetry: competitive equilibrium requires only local information—individual preferences and market prices—regardless of market size, whereas reaching the core entails increasingly comprehensive information about global cooperation potentials. This asymmetry raises fundamental questions about the theorem’s interpretation and practical applicability, particularly when we consider the cognitive limitations of real decision-makers.
To address this problem, one may consider of game theoretical models with incomplete information. Yet much of the existing literature on decision-making under incomplete information presupposes that players form probabilistic beliefs over a space of possible states. This assumption, deeply rooted in the Harsanyi’s doctrine, enables sophisticated modeling of strategic reasoning and communication. However, it also implicitly imposes a cognitively demanding structure: agents must possess subjective probability measures even when confronting profound uncertainty about the scope and nature of their ignorance.
This paper departs from that probabilistic framework to explore a different fundamental question: under what conditions can an agent justify a decision based solely on the information she actually possesses—without invoking probabilistic beliefs about unknown elements? Like Nasrudin in the epigraph, decision-makers often work with the information that is readily available rather than seeking what might be more relevant, either because they are unaware of what they do not know, or they recognize informational gaps but lack the motivation to fill them. We develop a model to capture this situation, where players possess partial, structural information about the situation, such as the values of specific coalitions, but neither quantify their ignorance through probability assignments nor engage in sophisticated reasoning about uncertainty.
We focus on the core, the canonical solution concept in cooperative games. The core, like other solution concepts, specifies a set of payoff vectors satisfying certain conditions, but the game formulation itself remains silent about what actions players can take and how these actions relate to achieving particular payoff vectors. To bridge this gap between the core and individual decision-making, we transform the cooperative game into a simple decision problem by asking each agent whether she would accept a proposed payoff vector. To give an answer, the agent must make logical inferences from the knowledge she possesses. [^2] The yes-or-no-to-a-proposal model bears similarity to Serrano and Vohra’s (2007). Yet our model does not require agents to engage in strategic sophistication or consider others’ behavior and beliefs. Our central theoretical concern is then to formalize what it means for information to be sufficient for “correct” acceptance with respect to the core. Rather than asking whether players can form optimal beliefs given their information, we ask whether the information they possess can logically support an unanimous acceptance of a proposed payoff vector if and only if the payoff vector is in the core.
This approach stands in marked contrast to existing models of cooperative games with incomplete information that rely on information partitions and belief hierarchies (Wilson, 1978; Serrano and Vohra, 2007; Liu, Mailath, Postlewaite, and Samuelson, 2014; Pomatto, 2022). [^3] Our model is in spirit closer to the model of unawareness (Heifetz, Meier, and Burkhard, 2013) and misspecification (Esponda and Puozo, 2016; Cerreia-Vioglio, Hansen, Maccheroni, Marinacci, 2025). Instead of modeling players’ strategic reasoning, we focus exclusively on one’s exogenous knowledge about the situation, that is, the potential chance of cooperation, without assuming that she has a theory of mind. We adopt this approach to identify the minimal informational skeleton underlying cooperation. By studying what each player can infer solely from her own information about the situation, we isolate the baseline conditions under which a payoff vector is acceptable for her. Such individual assessments form the base for cooperative communication, as cooperation typically begins when someone recognizes a potential Pareto improvement for a group of participants (Olson, 1965).
To define sufficiency of information, we need to determine the unit of information—that is, atomic sentences whose content concerns the decision-making situation. This requires a formal language that can precisely express statements about coalitional values. We modify the framework developed in Bonanno (2008, 2015) and syntactically encode the description of the situation. We model decision-making proof-theoretically by employing a Gentzen (1935a; 1935b)-style sequent calculus modified by Kaneko (2002); Kaneko and Suzuki (2002) to adapt to decision-theoretical analysis. [^4] The use of proof-theoretic methods to study strategic reasoning was initiated by Kaneko and Nagashima (1996, 1997). See Kaneko (2002) for a survey. Here, an agent’s reasoning process is formulated as a proof from ground assumptions (i.e., the information that she possesses) using classical inference rules. The decision criterion is formalized as a statement called Ci-acceptability: a payoff vector is Ci-acceptable for an agent if it cannot be logically refuted given the coalitional potentials known to her. Hence, whether to accept a proposed payoff vector $x$ is determined by whether the agent can prove that $x$ is Ci-acceptable based on her available information.
Our main result demonstrates that, under the minimal requirement that every coalition is known to at least one of its members, the set of payoff vectors that are unanimously Ci-acceptable coincides exactly with the core. Since obtaining and processing information burdens cognitive systems, this finding reinterprets the core not as a behavioral predictive solution but as a cognitive boundary. This reinterpretation illuminates a previously under-appreciated asymmetry in Debreu-Scarf theorem: while the mathematical convergence result remains valid, our analysis reveals that the informational requirements for justifying core outcomes may be prohibitively demanding in practice, suggesting that competitive equilibria—with their modest informational demands—may be more robust in the cognitive sense.
In general, a meaningful economic theory must grapple with cognitive limitations, recognizing that the most elegant theoretical results may be practically unattainable if they demand superhuman cognitive capabilities. By grounding cooperative game theory in logical derivability rather than probabilistic reasoning, this paper opens a new perspective on solution concepts in economics. Our approach builds bridges between individual reasoning, formal logic, and cooperative decision theory, offering a framework that, via information sufficiency, explores the cognitive limitations of real decision-makers while maintaining theoretical rigor. Further, the results provide insight into the minimal informational foundations required for cooperative behavior and suggest new directions for understanding how cooperation emerges under limited information.
The rest of the paper is organized as follows. Section 2 gives definitions of concepts in TU games. Section 3 gives a motivated introduction to formal language and their application in our model. Section 4 provides the main results. Section 5 discusses the cognitive inconsistency underlying Debreu-Scarf Theorem. All proofs are delegated to the Appendix.
2 TU Games and the Core: An Overview
This subsection provides an overview of cooperative games with transferable utility, as well as the concept of the core. For a more detailed treatment, see Kannai (1992) and Peleg and Sudhoelter (2007) (Chapters 2 and 3).
A cooperative games with transferable utility (TU game) is a pair $G=(N,v)$, where $N$ is the set of players and $v:2^{N}\rightarrow\mathbb{R}_{+}$ with $v(\emptyset)=0$ is the characteristic function. [^5] In this paper, $\mathbb{R}_{+}:=[0,\infty)$ is the set of non-negative real numbers, $\mathbb{N}:=\{1,2,$...\} is the set of positive integers, and $\mathbb{N}_{0}:=\mathbb{N}\cup\{0\}$ is the set of natural numbers. For simplicity, we restrict the codomain of characteristic functions to non-negative real numbers, which suffices for our analysis. The discussion extends naturally to characteristic functions with real-valued codomains. For each coalition $S\subseteq N$, $v(S)$ is interpreted as the maximal total payoff that can be obtained by $S$ (for example, through their collective action).
Given a cooperative game $G=(N,v)$, a payoff vector is a real-valued vector $x=(x_{i})_{i\in N}\in\mathbb{R}_{+}^{N}$, where each $x_{i}$ represents the amount of income received by player $i$. For each $S\subseteq N$, $x$ is feasible for $S$ if $\sum_{i\in S}x_{i}\leq v(S)$.
A central question in cooperative game theory is which payoff vectors should be considered “desirable” or acceptable given the structure of the game. One natural requirement is efficiency, or Pareto optimality, meaning that a payoff vector is a full distribution of the income earned by the comprehensive coalition $N$. Another compelling requirement is coalitional stability, which states that no subset of players should have an incentive to deviate and form a coalition that can secure a better outcome for its members. These two conditions together define the core, a canonical solution concept in cooperative game theory that captures efficient and stable (“fair”) payoff distributions.
Given $x,y\in\mathbb{R}^{N}$ and $S\subseteq N$, we say that $y$ dominates $x$ via $S$, denoted by $y$ domS $x,$ iff the following conditions hold: [^6] When only (i) holds, we say that $y$ weakly dominates $x$.
  • (i)
    $y_{i}\geq x_{i}$ for each $i\in S$;
  • (ii)
    $y_{i}>x_{i}$ for some $i\in S$.
Definition 1 (The core).
Given $G=(N,v).$ The core of $G$, denoted by $C(G)$, is the set of all payoff vectors $x=(x_{i})_{i\in N}$ that is feasible for $N$ and is not dominated by any payoff vector via any coalition. Formally, $x\in C(G)$ if and only if the following conditions are satisfied
  • Effeciency: $\Sigma_{i\in N}x_{i}=v(N)$;
[…]
Each $x\in C(G)$ is called a core payoff vector.
The core be defined alternatively as follows: $x\in C(G)$ if and only if the following two condition is satisfied:
  • Effeciency: $\Sigma_{i\in N}x_{i}=v(N)$;
  • Coalitional stability: For each $S\subseteq N$, $\Sigma_{i\in S}x_{i}\geq v(S)$.
Indeed, one can see that a payoff vector is coalitionally stable if and only if it is not dominated via any coalition $S$ by any payoff vector $y$ feasible for $S$.
Not every TU game has a non-empty core. Bondareva (1963) and Shapley (1967) characterize the non-emptiness of the core in terms of balancedness.
Due to the finitistic nature of the formal language—where formulas are of finite length and proofs proceed in finitely many steps—we assume that there exists an upper bound $M\in\mathbb{N}$ such that $v(S)\leq M$ for all coalitions $S$ in every TU game $G=(N,v)$ considered in this paper. [^7] This assumption is both technically convenient and realistic. In practice, $M$ could be the (rounded-up) global GDP. Similar boundedness assumptions are standard in the game theory literature (e.g., Shapley (1977)) and prove particularly valuable in competitive equilibrium analysis, as we shall see. Also, we focus on payoff vectors whose coordinates are natural numbers. Formally, let $G=(v,N)$ with $v(S)\in\mathbb{N}_{0}$. We let
$\textsf{PV}(G)=\{x=(x_{i})_{i\in N}\in[0,M]^{N}:\text{for each }i\in N,x_{i}\in\mathbb{N}_{0}\}$
We restrict discussions only on payoff vectors in $\textsf{PV}(G)$ to ensure finite divisibility (as payoffs in practice have smallest units, like cents in euro). An extension to rational coordinates will be discussed in Section 5.
3 Describing decision-making in a Formal Language
In our model, given a TU-game $G=(N,v)$, each agent $i\in N$ faces a decision-making situation in which she must either accept or reject a proposed payoff vector $x\in\mathbb{N}^{N}$. To make this decision, the agent must adopt a criterion for evaluation and determine—based on the information she possesses—whether the proposed payoff vector satisfies that criterion. This process involves drawing logical inferences from the agent’s knowledge. To formally analyze such reasoning, we require a syntactic framework that captures both the agent’s information and the structure of inference. This leads us to encode the game and the decision-making in a formal language.
Intuitively, a formal language has the following components: [^8] For a thorough introduction formal languages, inference rules, and proofs, see…. Though typically not part of the standard curriculum for economists, some game theorists use models built on formal logical language to study strategic reasoning in games. See, for example, Battigalli and Bonanno (1999), Arieli and Aumann (2015).
  • “Alphabet”. A (countable) set of atomic symbols, which are the indivisible units used to construct the expressions of a formal language.
  • “Sentences”. The set of “legitimate”, i.e, well-formed, strings of atomic symbols constructed according to the formation rules of the language.
  • “Ground assumptions”. A set of sentences that are assumed to be true without proof and serve as the basis for all inferences within the system.
  • “Reasoning rules”. A set of rules that determine the valid steps of reasoning, specifying how new sentences (conclusions) can be derived from existing ones (premises).
These four components determine what an agent can perceive and describe, how she reasons (“proofs”), and what conclusions she can reach (i.e., what she can “prove”). We first need to specify a formal language capable of capturing an agent’s reasoning in decision-making for yes-no problems within TU games. This section gives formal definitions of those concepts (Sections 3.1 for the “alphabet” and “sentences”, and Section 3.3 for the “ground assumptions”, “reasoning rules”, and “proofs”), and show how to apply them to the decision-making in the specific context of TU-games (Sections 3.2 and 3.4). Section 3.5 presents an example that illustrates how the various building blocks can be integrated to model an agent’s reasoning in the context of an aforementioned yes-no decision problem.
3.1 Formal Language in a Nutshell I: Alphabet and Sentences
To ground our discussion, we begin with the abstract definitions of alphabet and sentences—the foundational elements of any formal language. These definitions may appear austere at first glance, but their structure reflects deeper conceptual roles. In Section 3.2 we will explore their concrete interpretations and the intuitions that motivate their particular structure in our context.
“Alphabet”.
The “alphabet” is composed by the following symbols, which are the smallest unit of the language.
  • Atomic formulae: $\mathbf{p}_{0},\mathbf{p}_{1},\mathbf{p}_{2},$.... The set of atomic formulae is denoted by At.
  • Logical connectives: $\lnot$ (not), $\wedge$ (and), $\vee$ (or), $\rightarrow$ (implies); [^9] Many works, including Gentzen’s seminal papers (1935a; 1935b), use $\supset$ for implication. We follow textbooks such as Ebbinghause, Flum, and Thomas (2021) in using $\rightarrow$ as we believe this symbol is more intuitive for economists and game theorists.
  • Auxiliary symbols: parentheses: $($, );); braces: $\{$, };\}; commas: , .
“Sentences”.
Based on the symbols above, we define “legitimate” sequences of symbols in the alphabet, called formulae in the logic literature. A formula is obtained by a finite number of applications of the following three operations:
  • F0.
    Each atomic formula is a formula;
  • F1.
    If $A$ and $C$ are formulae, so are $\lnot A$ and $A\rightarrow C$;
  • F2.
    If $A_{0},A_{1},...,A_{m}$ are formulae, so are $\wedge_{k=1}^{m}A_{k}$ and $\vee_{k=0}^{m}A_{k}$;
We use F to denote the set of all formulae.
Here, one can see that our notion of “alphabet” differs significantly from its everyday usage: we are not referring to the familiar letters of a natural language like the twenty-six characters of the Latin script used in English. Instead, the alphabet consists of the minimal meaningful units within our formal system: the atomic propositions that serve as indivisible building blocks of logical reasoning. These atomic formulae represent basic statements that cannot be decomposed further. From these atomic components, we construct more complex propositions through systematic combination. By applying negation to a proposition, or by connecting two propositions with logical connectives such as “and”, “or”, or “implies”, we generate new, compound propositions. The complete set of all well-formed formulae in our language emerges through the finite, iterative application of the formation rules F0-F2, ensuring that every valid expression can be constructed through a finite sequence of rule applications from our basic alphabet.
So far, all symbols have been used in an abstract sense within the general definitions. In the following subsection, we will specify their concrete meaning in our particular context.
3.2 Alphabet and sentences for decision-making in TU games
Having introduced the general definition of alphabet and sentences, we now transform the abstract structure into a meaningful representation of the decision-making problem in our context. Recall that in our model, given a TU-game $G=(N,v)$, each agent $i\in N$ faces a decision-making situation in which she must either accept or reject a proposed payoff vector $x\in\mathbb{N}^{N}$. Our first task is to specify the alphabet and sentences that encode the relevant information for an agent’s decision-making.
Given $G=(N,v)$ with $v(S)\in\mathbb{N}$ for each $S\subseteq N$. Recall that $M\in\mathbb{N}$ is an upper bound of the payoffs that could be achieved by coalitions, i,e., $v(S)\leq M$ for each $S\subseteq N$, and we restrict our discussion on payoff vectors in which each coordinate is a non-negative integer and no larger than $M$, i.e., vectors in $\textsf{PV}(G)$. We define the set of atomic formulae $\textsf{At}(G):=\textsf{At}_{\text{fv}}(G)\cup\textsf{At}_{\text{com}}(G)$, where
$\textsf{At}_{\text{fv}}(G)$ $:=$ $\{x^{S}:x\in\textsf{PV}(G)\text{ and }S\subseteq N\},$
$\textsf{At}_{\text{wd}}(G)$ $:=$ $\{x\geq_{S}y:x,y\in\textsf{PV}(G)\text{ and }S\subseteq N\}.$
[…]
Sentences are constructed by a finite number of applications of formation rules F0 – F2 from $\textsf{At}(G)$ and the logical connectives and auxiliary symbols given in Section 3.1. We use $\textsf{F}(G)$ to denote the set of all sentences.
Here, one should distinguish between propositional symbols and their truth values. An atomic symbol $x^{S}\in\textsf{At}_{\text{fv}}(G)$ does not necessarily mean that the proposition it represents (i.e., “coalition $S$ can achieve payoff $x_{i}$ for each $i\in S$ through collective action”) is true. Similarly, $x\geq_{S}y\in\textsf{At}_{\text{wd}}(G)$ does not necessarily mean that $x$ weakly dominates $y$ via coalition $S$. These symbols merely represent propositions that agents can understand, use to construct more complex sentences, and reason about logically.
To simplify notation, in the following, we use $x>_{S}y$ as an abbreviation for the formula $x\geq_{S}y\wedge\lnot(y\geq_{S}x)$.
Example.
Consider a $2$-person TU game $G=(N,v)$ where $v(\{1\})=v(\{2\})=10$ and $v(\{1,2\})=30$. One can see that, for example, $(16,14)^{\{1\}}$ is an atomic formula in $\textsf{At}_{\text{fv}}(G)$, which is interpreted as “by staying alone, player $1$ can achieve payoff 16”. Note that the interpretation of $(16,14)^{\{1\}}$ is different from that of $(16,14)^{\{2\}}$, even as symbols they share the same payoff vector $(16,14)$: $(16,14)^{\{2\}}$ means that “by staying alone, player $2$ can achieve payoff 14”, which has nothing to do with player 1 and her payoff. Also, as an instance, one can see that $(8,7)\geq_{\{1\}}(9,2)$ and $(8,7)\geq_{\{2\}}(9,2)$ are two distinct atomic formulae in $\textsf{At}_{\text{wd}}(G)$; for each $i=1,2$, $(8,7)\geq_{\{i\}}(9,2)$ means that payoff vector $(8,7)$ weakly dominates the payoff vector $(9,2)$ via coaltion $\{i\}$, or, equivalently, player $1$ can reach payoff $9$ by staying alone, which is better than $8$, and player $2$ can reach payoff $2$ by staying alone, which is better than $7$.
Note that, as we mentioned before, not all the formulae are true in the context: clearly, $(16,14)^{\{i\}}$ are false for both $i=1,2$, because $v(\{1\})=v(\{2\})=10$ and neither 16 or 14 is reachable by a player’s staying alone. Also, $(8,7)\geq_{\{1\}}(9,2)$ is also false, because $8<7$. Nevertheless, all are atomic symbols: agents understand them, and, when making a decision to the yes-no question, an agent may have to determine their truth value by starting from some ground assumptions and going through some inference rules, which will be introduced in the next two subsections.
3.3 Formal Language in a Nutshell II: Axioms, Inference Rules, and Proof
So far, we have laid out the static backbone of our logical language—its alphabet of symbols and the well-formed sentences (or formulae) that can be constructed from them. Think of this as having assembled the “raw materials” of our logical language. However, a language alone does not constitute a reasoning system. Just as knowing the words and grammar of English doesn’t automatically tell us which statements are true or which arguments are valid, our logical alphabet and sentence formation rules do not yet provide us with any means of distinguishing valid reasoning from invalid reasoning. What remains is the dynamic machinery that brings this language to life: a systematic method for generating valid sentences through appropriate logical inferences.
In such a method, we begin by stipulating certain axioms—statements we accept without proof—and then apply inference rules to derive new, valid sentences from those already established. Among the various approaches —including the well-known Hilbert-style systems and Gentzen’s natural deduction—we will employ Gentzen’s sequent calculus, a framework that has proven particularly powerful for understanding the deep structure of logical reasoning and has become indispensable in modern proof theory.
This subsection provide a brief introduction to the sequent calculus. A sequent is an expression in the form
$\Gamma\Rightarrow\Theta$
Here $\Gamma$ and $\Theta$ are finite (probably empty) sets of formulae (i.e., sentences); $\Gamma$ is called the antecedent and $\Theta$ the succedent. [^11] Many works, including Gentzen’s seminal papers (1935a; 1935b), use $\rightarrow$ instead of $\Rightarrow$. Here, to avoid confusing with implication in formulae, we follow textbooks such as Mancosu, Galvan, and Zach (2021) and use $\Rightarrow$.
[…]
In the following, we use Greek letters for set of formulae and Latin letters for formulae. Following the convention in the literature, given a set $\Gamma$ of formulae and a formula $A$, we write $A,\Gamma$ or $\Gamma,A$ instead of $\Gamma\cup\{A\}$ at the antecedent or the succedent to emphasize that a formula $A$ is added to the set $\Gamma$.
In sequent calculus, a proof (derivation) is defined by a tree of sequents, where each terminal node is an axiom, and predecessor nodes (sequents) are connected with their successor node (sequent) by an inference rule. At the root is the sentence proved by the proof. A sequent $\Gamma\Rightarrow\Theta$ is called provable, denoted by $\vdash\Gamma\Rightarrow\Theta$, when it can be proved by a proof.
We start from the inference rules. Following the Gentzen’s LK sequent calculus system modified by Kaneko (2002), they are:
Structural rules:
$\frac{\Gamma\Rightarrow\Theta}{\Delta,\Gamma\Rightarrow\Theta,\Lambda}\text{(Weakening) \ \ \ \ \ \ \ \ \ \ }\frac{\Gamma\Rightarrow\Theta,A\ \ \ \ A,\Delta\Rightarrow\Lambda}{\Delta,\Gamma\Rightarrow\Theta,\Lambda}\text{(Cut)}$
Operational rules:
$\frac{\Gamma\Rightarrow\Theta,A}{\lnot A,\Gamma\Rightarrow\Theta}(\lnot\text{-Left) \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ }\frac{A,\Gamma\Rightarrow\Theta}{\Gamma\Rightarrow\Theta,\lnot A}\text{(}\lnot\text{-Rigt}\text{)}$
$\frac{\Gamma\Rightarrow\Theta,A\ \ \ \ B,\Gamma\Rightarrow\Theta}{A\rightarrow B,\Gamma\Rightarrow\Theta}(\rightarrow\text{-Lef) \ \ \ \ \ \ \ \ \ \ \ }\frac{A,\Gamma\Rightarrow\Theta,B}{\Gamma\Rightarrow\Theta,A\rightarrow B}\text{(}\rightarrow\text{-Right}\text{)}$
$\frac{A,\Gamma\Rightarrow\Theta}{\wedge\Phi,\Gamma\Rightarrow\Theta}(\wedge\text{-Left}\text{) where }A\in\Phi\text{ \ \ \ \ \ \ \ \ \ \ \ \ }\frac{\{\Gamma\Rightarrow\Theta,A:A\in\Phi\}}{\Gamma\Rightarrow\Theta,\wedge\Phi}\text{(}\wedge\text{-Right}\text{)}$
[…]
Here, each Greek letter, such as $\Gamma$, $\Theta$, $\Delta$, $\Delta$, and $\Phi$, represents a finite (probably empty) set of formulae, and each Latin letter, such as $A$ and $B$, represent a single formula. The symbol $\wedge\Phi$ ($\vee\Phi$) represents a formula formed by connecting all the elements in $\Phi$ by “and” (“or”); formally, given $\Phi=\{A_{1},...,A_{m}\}$, $\wedge\Phi:=A_{1}\wedge...\wedge A_{m}$ and $\vee\Phi:=A_{1}\vee...\vee A_{m}$.
Each inference rule is composed by a line in the middle, a group of sequents above the line, called upper sequent, and one sequent below the line, called lower sequent; it indicates that the lower sequent, as the conclusion, can be inferred from the upper sequents as hypotheses. The twelve inference rules are divided into two groups. The structural rules deal with “diluting” and “condensing” the structure of a sequent. The operational rules deal with logical operators. [^13] The Cut rule is a special inference rule that differs from others in an important way: in all other rules, every formula occurring in the premise(s) also appears in the conclusion, either directly or as a subformula. The Cut rule, by contrast, literally “cuts” a formula out of the derivation. As a result, when the Cut rule is used in a proof, it becomes difficult to trace back from the conclusion to the necessary premises. Gentzen (1935a) showed that any formula provable using the Cut rule can also be proven without it—a result known as the Cut-elimination theorem, or Gentzen’s Hauptsatz, which is foundational in modern proof theory. Nevertheless, following common practice, we retain the Cut rule here, as it can significantly simplify proofs. For a detailed discussion, see Boolos (1984). An inference rule is composed by two parts divided by a horizontal line in the middle. Each sequent above the line is called a premise of the rule, and the sequent below the line the conclusion of the rule. Each inference rule specifies the derivation of a conclusion from premises.
[…]
The inference rules of sequent calculus are significant because they provide a systematic, symmetric framework where logical operations are broken down into their most basic proof-theoretic components, making the structure of logical reasoning transparent and enabling powerful meta-theoretical results like cut-elimination (see footnote 13), providing a foundation for understanding computation and proof search.
In addition to using inference rules to connect nodes, a proof tree must begin with axioms. A formal language typically has two types of axioms: logical axioms and non-logical axioms.The KD system has simple logical axioms, namely, all tautologies:
  • Logical axiom schema: $A\Rightarrow A$
It is called axiom scheme because it provides a scheme instead of listing all axioms: here, $A$ could be replaced by any formula. Non-logical axioms are the ground assumptions specific to a particular context. In the following subsection, we present the non-logical axioms in our model and demonstrate how to formalize an agent’s reasoning through formal proofs.
3.4 Reasoning and Decision-making with respect to TU games
To capture how agents reason about our yes-no decision problem, we need to equip them with the ability to make basic numerical comparisons. Intuitively, agents should recognize when one outcome vector is at least as good as another with respect to a coalition — that is, when every component is at least as good. We formalize this intuition through the following non-logical axioms:
Non-logical axioms:
  • $\text{ \ \ \ }\Rightarrow y\geq_{S}x$ for all $y\geq_{S}x\in\mathsf{At}_{\text{wd}}$ with $y_{i}\geq x_{i}$ for each $i\in S$;
  • $\text{ \ \ \ }\Rightarrow\lnot(y\geq_{S}x)$ for all $y\geq_{S}x\in\mathsf{At}_{\text{wd}}$ with $y_{i}<x_{i}$ for some $i\in S$.
Note that all those axioms are sequent with an empty antecedent, meaning we treat them as fundamental truths that require no additional justification. The first group of axioms establishes that $y\geq_{S}x$ with respect to coalition $S$ whenever each member $i\in S$ receives at least as much in $y$ as in $x$; the second axiom ensures that its negation, i.e., $\lnot(y\geq_{S}x)$ holds whenever any coalition member in $S$ is strictly worse off. For instance, in Example 3.1,  $\text{ \ \ \ }\Rightarrow(8,7)\geq_{\{2\}}(9,2)$ is an axiom, while  $\text{ \ \ \ }\Rightarrow(8,7)\geq_{\{1\}}(9,2)$ is not.
In summary, these axioms establish that $y\geq_{S}x$ holds if and only if $y_{i}\geq x_{i}$ for all $i\in S$. This gives our agents the essential capability to recognize unambiguous improvements—a cornerstone of rational decision-making.
3.5 An Example
We conclude this section by a detailed example that shows, step-by-step, how all these building blocks – alphabets, sentences, axioms, inference rules — are combined together to describe an agent’s decision-making.
Example.
Consider the $2$-person TU game in Example 3.1 where $v(\{1\})=v(\{2\})=10$ and $v(\{1,2\})=30.$ Note that the core of the $G$ is $\{(x,30-x):10\leq x\leq 20\}$. Here we take player 1’s viewpoint. Suppose that she faces a proposal in the core with integer coordinates, say $x=(14,16)$. How should we describe her reasoning to accept or reject the proposal? An intuitive idea coherent to cooperative game theory is that she rejects the proposal if and only if, based on the information she possesses, she could achieve a higher payoff either by working alone or cooperating with agent $2$. To capture this intuition, consider the following two sentences (i.e., formulae):
$\bigvee_{y\in\textsf{PV}(G)}\big(y^{\{1\}}\wedge(y>_{\{1\}}x)\big)$ (1)
$\bigvee_{y\in\textsf{PV}(G)}\Big(y^{\{1,2\}}\wedge(y\geq_{\{1,2\}}x)\wedge(y>_{\{1\}}x)\Big)$ (2)
Sentence (1) describes that agent $1$ can improve upon $x$ by herself, that is, there is a “considerable” payoff vector $y$ such that (i) $y$ is feasible for agent $1$ by working alone, that is, $y^{\{1\}}$, and (ii) what agent $1$ gets in $y$ is strictly better than what she gets in $x$, that is, $(y>_{\{1\}}x)$.
Sentence (2) describes that agent $1$ can improve upon $x$ by cooperating with agent 2, that is, there is a “considerable” payoff vector $y$ such that (i) $y$ is feasible if $1$ and $2$ cooperate, that is, $y^{\{1,2\}}$, (ii) they have no motivation to resist this cooperation because no one is worse off in $y$ than in $x$, that is, $y\geq_{\{1,2\}}x$, and (iii) agent $1$ has strict motivation to achieve this cooperation because what agent $1$ gets in $y$ is strictly better than what she gets in $x$, that is, $(y>_{\{1\}}x)$.
To summarize, agent $1$ should reject the proposal $x$ if and only if she could prove sentence (1) or (2). Equivalently, she accepts $x$ if and only if she could prove that both sentences are wrong, that is, she could prove “neither sentence (1) or sentence (2) holds”, which is
$\lnot\left(\bigvee_{y\in\textsf{PV}(G)}\big(y^{\{1\}}\wedge(y>_{\{1\}}x)\big)\vee\bigvee_{y\in\textsf{PV}(G)}\left(y^{\{1,2\}}\wedge(y\geq_{\{1,2\}}x)\wedge(y>_{\{1\}}x)\right)\right)$ (3)
We use $C_{1}(x)$ to denote the formula in (3). Now comes a crucial question: can agent 1 actually prove $C_{1}(x)$ using only the axioms we have introduced so far? Formally, is  $\text{ \ \ }\Rightarrow C_{1}(x)$ provable? At first glance, one might think the answer is yes—but a closer look reveals a fundamental gap.
The problem is this: to prove $C_{1}(x)$, the agent needs information about which outcomes are actually feasible (i.e., which $y^{S}$ statements are true or false). But our axioms only tell us about numerical comparisons—they say nothing about what outcomes are actually available to different coalitions. It is like having rules for comparing temperatures but no thermometer to take actual measurements.
More precisely, our current axioms cannot generate any statement of the form $y^{S}$ or $\lnot y^{S}$. Without this information, the agent cannot determine whether better alternatives exist, making it impossible to prove either $C_{1}(x)$ or its negation $\lnot_{1}C(x)$. The inference rules we have defined cannot conjure this missing information out of thin air. [^14] Formally, recall that every proof is a tree, where in each terminal node there a (logical or non-logical) axiom, and in the node the sentence that we want to prove. To show that $y^{S}$ or $\lnot y^{S}$, the formula should enter into the proof somewhere. However, none of the axioms includes any information in the form of $y^{S}$; the inference rule Weakening can introduce formula “out of blue”, but cannot force it into a conjunct.
Therefore, we cannot prove  $\text{ \ \ }\Rightarrow C_{1}(x)$ but need to provide the sequent with some antecedent, that is, agent $1$ needs some information to derive $C_{1}(x)$. Here, the relevant information would be about the payoff of cooperation, i.e., the characteristic function $v$. Therefore, we can suppose that agent 1 has information about the payoff generated by every coalition including her. This information could be formally described as a finite set of sentences $\hat{\Gamma}(\mathcal{N}_{1})$ as follows:
[…]
Here, $\hat{\Gamma}(\mathcal{N}_{1})$ provide “full” information about $v$ for agent $1$ to make decision. It is composed by four parts. The first and the third are “positive”: they specify what is reachable for agent 1 by working alone (the first line on 4) and by cooperating with agent 2 (the third line). The second and fourth are “passive”: they specify what is unreachable by working alone (the second line) and by cooperating with agent 2 (the fourth line).
Now we show what happens when we give agent 1 complete information. If agent 1 is equipped with full knowledge about all possible cooperation outcomes in $G$ — that is, the “full” information represented by $\hat{\Gamma}(\mathcal{N}_{1})$ —she can finally reduce some conclusion that axioms alone cannot help her to reduce.
In other words, armed with full information the payoff generated by any coalition includes her, she can prove $C_{1}(x)$: no feasible alternative payoff vector can give her a better payoff than 14, i.e., the payoff she receives under proposal $x$. Formally, this means that $\hat{\Gamma}(\mathcal{N}_{1})\Rightarrow C_{1}(x)$ is provable.
To show this, an intuitive reasoning is that, given any payoff vector $y\in\textsf{PV}(G)$, either, by the information codified in $\hat{\Gamma}(\mathcal{N}_{1})$, the agent derives that it is not feasible, or by numerical comparison (that is, by non-logical axioms), the agent implies that it is not an improvement (or, even if it is an improvement for $1$, agent 2 would not have any motivation to cooperate). To see how this intuition can be captured by a formal proof, consider, for example, $y\in\textsf{PV}(G)$ with $y^{\{1\}}\not\in\hat{\Gamma}(\mathcal{N}_{1})$, that is, $\lnot y^{\{1\}}\in\hat{\Gamma}(\mathcal{N}_{1})$. We have the following proof:
$\frac{\frac{\frac{\lnot y^{\{1\}}\Rightarrow\lnot y^{\{1\}}}{\hat{\Gamma}(\mathcal{N}_{1})\Rightarrow\lnot y^{\{1\}}}(\text{Weakening})\text{ \ \ \ }\frac{y^{\{1\}}\Rightarrow y^{\{1\}}}{\lnot y{\{1\}},y^{\{1\}}\Rightarrow\text{ \ }}(\lnot\text{-Left})}{y^{\{1\}},\hat{\Gamma}(\mathcal{N}_{1})\Rightarrow\text{ \ \ \ }}(\text{Cut})}{y^{\{1\}}\wedge(y>_{\{1\}}x),\hat{\Gamma}(\mathcal{N}_{1})\Rightarrow\text{ \ \ \ }}(\wedge\text{-Left})$ (5)
The proof starts with two terminal nodes. The left branch begins with a logical axiom (a tautology) $\lnot y^{\{1\}}\Rightarrow\lnot y^{\{1\}}$; through Weakening, this becomes $\hat{\Gamma}(\mathcal{N}_{1})\Rightarrow\lnot y^{{1}}$. This captures the essential intuition: given the information the agent possesses (encoded in $\hat{\Gamma}(\mathcal{N}_{1})$), she derives that $y$ is not feasible for her.
[…]
It would be helpful if we consider the complementary case, namely $y\in\textsf{PV}(G)$ with $y^{\{1\}}\in\hat{\Gamma}(\mathcal{N}_{1})$ (that is, by definition, $y_{1}\geq v(\{1\})=10<14=x_{1}$). Here, though $y$ is feasible, agent should reject $y$ by the comparison ability given non-logical axioms. The formal proof is as follows:
 $\frac{\frac{\frac{\text{ \ \ }\Rightarrow x\geq_{\{1\}}y}{\lnot(x\geq_{\{1\}}y)\Rightarrow\text{ \ \ }}(\lnot\text{-Left})}{y^{\{1\}}\wedge(y\geq_{\{1\}}x)\wedge\lnot(x\geq_{\{1\}}y)\Rightarrow\text{ \ \ }}(\wedge\text{-Left})}{y^{\{1\}}\wedge(y>_{\{1\}}x),\hat{\Gamma}(\mathcal{N}_{1})\Rightarrow\text{ \ \ \ }}(\text{Weakening})$ (6)
The proof starts from a non-logical axiom  $\text{ \ \ }\Rightarrow x\geq_{\{1\}}y$, which accurately captures the aforementioned intuition. The logical machinery then generates the desirable $y^{\{1\}}\wedge(y>_{\{1\}}x),\hat{\Gamma}(\mathcal{N}_{1})\Rightarrow\text{ \ \ \ }$: it has been demonstrated that an agent equipped with basic numerical comparison ability cannot consider $y$ a feasible improvement on $x$ if she works alone, even though she has information about the cooperation payoffs.
[…]
$\left(\bigvee_{y\in\textsf{PV}(G)}\big(y^{\{1\}}\wedge(y>_{\{1\}}x)\big)\vee\bigvee_{y\in\textsf{PV}(G)}\left(y^{\{1,2\}}\wedge(y\geq_{\{1,2\}}x)\wedge(y>_{\{1\}}x)\right)\right),\hat{\Gamma}(\mathcal{N}_{1})\Rightarrow$
By applying $\lnot$-Right rule, we obtain $\hat{\Gamma}(\mathcal{N}_{1})\Rightarrow C_{1}(x)$. Here, we have shown that $\hat{\Gamma}(\mathcal{N}_{1})\Rightarrow C_{1}(x)$ is provable, that is, equipped with basic reasoning ability, through information $\hat{\Gamma}(\mathcal{N}_{1})$ agent 1 implies that she cannot improve her payoff in $x$ by working alone or cooperating with agent 2. Hence she should accept $x$.
The logic becomes even clearer when we consider payoff vectors that can be improved upon. Take, for instance, $\tilde{x}=(11,12)$—here we can formally establish that $\vdash\hat{\Gamma}(\mathcal{N}_{1})\Rightarrow\lnot C_{1}(x)$, meaning agent 1’s information logically compels her to reject this allocation. This demonstrates how (basic) reasoning ability, when properly informed, naturally filters out inefficient outcomes.
But this raises deeper questions about the information structure itself. What happens when agents operate under incomplete information? Suppose agent 1 knows her $v({\{1\}})$ but does not know the grand coalition’s worth $v({\{1,2\}})$ (she might even unaware of the potential of cooperation). Can she still make acceptance decisions?
Further, even knowing $v({\{1\}})$ and $v({\{1,2\}})$, agent $1$ will not reject $x^{\prime}=(21,9)$: based on her information, she still could derive $C_{1}(x^{\prime})$. Yet agent 2 will reject it as far as agent 2 possesses information about $v(\{2\})$. Since the core, as all solution concepts, concerns payoff distributions that would “satisfy” all participants, it is reasonable to take into account not one agent but all agents’ decision to the yes-no question. In that case, how does the strategic landscape change?
Ultimately, all these individual decision problems converge on a fundamental question for cooperative game theory: if a payoff vector is implemented only when all agents accept it, what informational conditions ensure that the accepted outcomes coincide precisely with the core? In other words, how much information must be distributed among the players, and in what configuration, to guarantee that individual reasoning ability aggregates into collective efficiency? The next section formalizes this connection between information, individual logic, and core stability.
4 Information Sufficiency for Unanimous Acceptance of Core Payoff Vectors
Given a TU game $G=(N,v)$ and a proposed payoff vector $x\in\textsf{PV}(G)$, how should an agent decide whether to accept or reject $x$? Example 3.2 suggests a natural approach: the (private) information possessed by an agent and her decision criterion are encoded as formal sentences, and her reasoning process is be described by logical proofs; her acceptance-refusal choice depends upon whether the criterion is provable given her information and her reasoning ability.
This approach transforms game-theoretic questions into information- and proof-theoretical ones. Rather than asking “what makes the core stable?” we can ask “what information enables agents to recognize core allocations”. Example 3.2 demonstrated this approach for two-player games, but the real challenge lies in scaling up: as coalitions proliferate and information becomes distributed across multiple agents, the logical landscape grows exponentially more complex.
This section generalizes Example 3.2’s approach to address three fundamental questions:
  • Q1.
    How can we formulate a general decision criterion that captures the core’s underlying intuition across arbitrary games?
  • Q2.
    What informational requirements must be satisfied for an individual agent to reliably accept core payoffs vectors?
  • Q3.
    When all agents employ this criterion, what distribution of information ensures that unanimous acceptance occurs if and only if the payoff vector lies in the core?
These questions bridge individual cognition and collective outcomes, revealing how the core emerges from the interplay between information and logical reasoning.
We first answer Q1. Let $G=(N,v)$ with $v(S)\in\mathbb{N}$ for each $S\subseteq N$. For each $i\in N$, we define $\mathcal{N}_{i}:=\{S\subseteq N:i\in S\}$, i.e., the set of all coalitions that include $i$. Given a payoff vecyor $x\in\textsf{PV}(G)$, consider the following formula:
$\lnot\left(\bigvee\limits_{S\in\mathcal{N}_{i}}\bigvee\limits_{y\in\mathsf{PV}(G)}(y^{S}\wedge(y\geq_{S}x)\wedge(y>_{\{i\}}x)\right)$ (CORE-CRIT)
Comparison with the formula in (3) shows that the formula in (CORE-CRIT) is an extension of the former. It means that agent $i$ cannot find any coalition $S$ containing herself and any payoff vector $y\in\textsf{PV}(G)$ such that
  • (i)
    $y$ is feasible for coalition $S$, described by $y^{S}$;
  • (ii)
    In $y$, the payoff for everyone in $S$ is as good as in $x$, described by $y\geq_{S}x$;
  • (iii)
    Agent $i$ gets a strictly higher payoff in $y$ than in $x$, described by $y>_{\{i\}}x$.
We denote the formula in (CORE-CRIT) as $C_{i}(x)$. This expression elegantly captures how each agent must evaluate a payoff vector: she cannot join any coalition to strictly improve her payoff. In essence, $C_{i}(x)$ formalizes the individualized logic underlying core stability—it translates the question “is it a core payoff vector?” into “can I prove that no profitable deviation exists?”
By adopting $C_{i}(x)$ as agent $i$’s decision criterion, we achieve a fundamental transformation: the core’s stability properties, traditionally understood in terms of coalitional incentives, become questions of logical provability for individual agents. Each player’s acceptance decision now depends not on what is true about the game, but on what she can demonstrate given her available information. Given agent $i$’s information $\Gamma_{i}$, which is a subset of sentences, a payoff vector $x$ is $\textsf{C}_{i}$-acceptable under $\Gamma_{i}$ if and only if $\Gamma_{i}\rightarrow C_{i}(x)$ is provable.
However, a crucial complication emerges: agents may face undecidability. As mentioned in Example 3.2, there are payoff vectors $x$ where an agent’s information $\Gamma_{i}$ leaves the criterion $C_{i}(x)$ in logical limbo—neither provable nor refutable. In such cases, the agent cannot derive either $C_{i}(x)$ or its negation $\lnot C_{i}(x)$ from her available information, leaving her without a clear accept-or-reject decision.
This undecidability problem, fundamental to both logic and computer science, poses a basic question for our framework: what informational conditions guarantee that agents can always reach determinate conclusions? Formally, when can we ensure that for any payoff vector $x$, either $\Gamma_{i}\vdash C_{i}(x)$ or $\Gamma_{i}\vdash\lnot C_{i}(x)$ holds? Understanding these conditions will reveal the minimal information requirements for coherent decision-making in TU games.
As Example 3.2 suggests, in (CORE-CRIT), the decidability relies up $y^{S}$, because both $y\geq_{S}$ and $y>_{\{i\}}x$ depend only on the numerical comparison abilities encoded in the non-logical axioms. Hence, effective information should only contain $y^{S}\in\textsf{At}_{\text{fv}}$ or their negations. Here, for notational convenience, we slightly abuse notation by letting $\lnot\mathsf{At}_{\text{fv}}:=\{\lnot y^{S}:y^{S}\in\mathsf{At}_{\text{fv}}\}$. The following proposition provides a sufficient condition for decidability.
Decidability and consistency Proposition.
If $\Gamma_{i}\subseteq\mathsf{At}_{\emph{fv}}\cup\lnot\mathsf{At}_{\emph{fv}}$ and satisfies the following condition:
  • (D)
    For each $y^{S}\in\mathsf{At}_{\emph{fv}}$, either $y^{S}\in\Gamma_{i}$ or $\lnot y^{S}\in\Gamma_{i}$ but not both,
then for each $x\in\mathsf{PV}(G)$, either $\vdash\Gamma_{i}\rightarrow C_{i}(x)$ or $\vdash\Gamma_{i}\rightarrow\lnot C_{i}(x)$, but not both.
Condition (D) ensures that $\Gamma_{i}$ provides a definite judgment—affirmative or negative—on every atomic statement about feasibility, i.e., for each $y^{S}\in\mathsf{At}_{\text{fv}}$, $\Gamma_{i}$ includes either $y^{S}$ or its negation, but not both. In other words, $\Gamma_{i}$ is a logically maximal consistent set over $\mathsf{At}_{\text{fv}}$. This “completeness” ensures that for any $x\in\mathsf{PV}(G)$, the agent can determine whether $x$ is $\mathsf{C}_{i}$-acceptable or not. That is, the agent avoids ambiguity or contradiction and makes consistent judgments.
Note, however, that for particular games and vectors, not all formulas in a maximal set $\Gamma_{i}$ are required to decide acceptability. For example, in Example 3.2, $(14,16)$ is $\mathsf{C}_{1}$-acceptable with respect to $\hat{\Gamma}(\mathcal{N}_{1})$, while $\hat{\Gamma}(\mathcal{N}_{1})$ does not satisfy (D). Still, Proposition 4.3 holds at a general level: without knowing $v$, condition (D) guarantees that the agent’s reasoning about any $x$ is always decidable. If (D) fails, there may exist some characteristic function $v$ for which the $\mathsf{C}_{i}$-acceptability of some $x$ remains undecidable.
While condition (D) ensures logical maximal consistency, it is relatively loose from a contextual perspective: a set $\Gamma_{i}$ can satisfy (D) while still encoding information that is contextually unreasonable about coalition formation. For instance, it is possible for $\Gamma_{i}$, which satisfies (D), to include both $y^{S}$ and $\lnot\tilde{y}^{S}$ for some vectors $y,\tilde{y}$ such that $\sum_{j\in S}\tilde{y}_{j}<\sum{j\in S}y_{j}$. In such a case, $\Gamma_{i}$ affirms that $y$ is feasible for $S$ while denying the feasibility of $\tilde{y}$, even though $\tilde{y}$ requires fewer resources than $y$ and should therefore be at least as feasible. This illustrates that $\Gamma_{i}$ may satisfy (D) without conveying any meaningful or reasonable information about coalition values.
To address this, we may wish to impose context-sensitive restrictions on the content of $\Gamma_{i}$, ensuring that the information it encodes reasonably reflects the structure and potential of coalition formation. Formally, given $G=(N,v)$. For each $\mathcal{S}\subseteq 2^{N}$, we define
$\hat{\Gamma}^{G}(\mathcal{S}):=\{y^{S}:y^{S}\in\textsf{At}_{\text{fv}},S\in\mathcal{S}\text{, and }\sum_{j\in S}y_{j}\leq v(S)\}$
And we let
$\Gamma^{G}(\mathcal{S}):=\hat{\Gamma}(\mathcal{S})\cup\{\lnot y^{T}:y^{T}\in\textsf{At}_{\text{fv}}\setminus\hat{\Gamma}(\mathcal{S})\}$
Note that the superscript $G$ is important: for $G=(N,v)$ and $G^{\prime}=(N,v^{\prime})$ with $v\neq v^{\prime}$, it is possible that $\Gamma^{G}(\mathcal{S})\neq\Gamma^{G^{\prime}}(\mathcal{S})$. One can see easily that for each $G=(N,v)$ and each $\mathcal{S}\subseteq 2^{N}$, $\Gamma(\mathcal{S})$ is a subset of $\textsf{At}_{\text{fv}}\cup\lnot\textsf{At}_{\text{fv}}$ and satisfies condition (D).
Given a game $G=(N,v)$, the set $\Gamma^{G}(\mathcal{S})$ encodes the agent’s information about the feasibility of payoff vectors for coalitions in $\mathcal{S}$. Specifically, for each $S\in\mathcal{S}$, any formula $y^{S}$ satisfying $\sum_{j\in S}y_{j}\leq v(S)$ is included in $\Gamma^{G}(\mathcal{S})$, affirming its feasibility for $S$. This reflects a situation where the agent has complete and accurate information about the capabilities of the coalitions in $\mathcal{S}$.
For all coalitions outside of $\mathcal{S}$, every possible payoff vector $y^{T}$ is rejected, i.e., $\lnot y^{T}\in\Gamma^{G}(\mathcal{S})$. This captures a strong modeling assumption: when an agent lacks information about a coalition—whether due to unawareness, ignorance, or uncertainty—they systematically deny all payoff possibilities for that coalition. In other words, $\Gamma^{G}(\mathcal{S})$ packages epistemic limitations (such as unawareness or insufficient information) into a syntactic simplification: total denial of feasibility for unknown coalitions. [^15] While this may seem extreme, as shown in the behavior of Nasrudin in the epitaph, this is not abnormal. Neither it is without precedent in the literature. A similar idea appears in decision-theoretic treatments of ambiguity and model misspecification where states outside an agent’s belief model are assigned utility $\infty$ (see Gilboa and Schmeidler, 1989). By relaxing our restriction and introducing ideas from fuzzy logic, one may bridge our model with the literature on incomplete information in cooperative games. Further research is needed in this direction. In what follows, we restrict our attention to information sets of the form $\Gamma^{G}(\mathcal{S})$ for some $\mathcal{S}\subseteq 2^{N}$.
Based on this assumption, the next result answers Q2, that is, what informational requirements must be satisfied for an individual agent to reliably accept core payoffs vectors.
Core payoff vectors are always $\textsf{C}_{i}$-acceptable Proposition.
Given $G=(N,v)$ and let $x\in\mathsf{PV}(G)$. If $x\in C(G)$, then $\vdash\Gamma^{G}(\mathcal{S})\Rightarrow C_{i}(x)$ for each $\mathcal{S}\subseteq 2^{N}$.
Proposition 4.4 states that any payoff vector in the core of a game is $\mathsf{C}_{i}$-acceptable regardless of which coalitions $\mathcal{S}$ the agent has information about. The intuition is straightforward: if an agent has no information (i.e., $\mathcal{S}=\emptyset$), she cannot refute any vector—there is no known feasible alternative to compare. If she possesses some information, then, because the vector lies in the core, none of the coalitions she knows about can offer a deviation. Thus, core payoffs remain unrefutable from any informational standpoint.
However, this result is partial in a crucial sense: it tells us that core vectors are always acceptable, but it does not characterize when non-core vectors become unacceptable. Naturally, this requires more information—an agent must know enough to witness a feasible deviation. The following observation clarifies which kind of information matters. Specifically, if an agent receives additional information about coalitions that do not contain her, this does not change her ability to judge whether a payoff vector is not $\mathsf{C}_{i}$-acceptable. That is, knowing about coalitions she does not belong to is irrelevant to her own refutation process. Only information about coalitions including the agent can affect her judgment of acceptability.
Observation 1 (Information about irrelevant coalitions is useless).
Given $G=(N,v)$ and let $\mathcal{S},\mathcal{T}\subseteq 2^{N}$ with $\mathcal{S}\cap\mathcal{T}=\emptyset$. If $\mathcal{T}\cap\mathcal{N}_{i}=\emptyset$ (i.e., $i\not\in T$ for each $T\in\mathcal{T}$), then $\vdash\Gamma^{G}(\mathcal{S})\Rightarrow C_{i}(x)$ if and only if $\vdash\Gamma^{G}(\mathcal{S}\cup\mathcal{T})\Rightarrow C_{i}(x)$.
Observation 1 shows that information about coalitions not involving agent $i$ is irrelevant to her judgment about $\mathsf{C}_{i}$-acceptability. A direct consequence is that even if a payoff vector lies outside the core, agent $i$ may still find it acceptable, provided she cannot improve her own payoff in it. For example, as illustrated in Example 3.2, the vector $(21,9)$ is not in the core because it violates agent 2’s individual rationality. However, agent 1 does not refute it — even when she knows all coalitions that include her — since she sees no feasible deviation for herself. By contrast, agent 2 would refute it, assuming she knows v(2)v({2}).
This underscores a crucial point: $\mathsf{C}_{i}$-acceptability is inherently agent-relative. Hence, to evaluate a vector, we need to take into account of every agent’s answer to the proposed vector and consider the unanimous acceptability. The next result offers a preliminary insight into this: if each agent knows all coalitions that include her, then only payoff vectors in the core are accepted by all. That is, full self-relevant information across all agents recovers the core as the set of unanimously acceptable payoff vectors.
Observation 2 (Full information characterizes the core).
For each $x\in\mathbb{N}^{N}$, $x\in C(G)$ if and only if $\vdash\Gamma^{G}(\mathcal{N}_{i})\Rightarrow C_{i}(x)$ for each $i\in N$.
Observation 2 is a corollary of Theorem 4.5 in the following, and we omit its proof. Since the size of $\mathcal{N}_{i}$ could be large even with a moderate size of $|S|$, one may wonder whether the unanimous acceptance only of the core vector could be achieved by smaller sets. Our main result provides a (partially) tight condition to the question.
It should be noted that this is a decentralization of knowledge in the sense that a player does not need to know all coalitions. On the other hand, when $N$ grows very large, this is demanding on a player’s cognitive ability since she has to know $2^{|N|-1}$ coalitions. Hence it is natural to wonder whether each player really needs to know so many coalitions for reaching an agreement only on core-payoff vectors. We have the following theorem, which characterize unanimous acceptance of only core payoffs in terms of the minimal knowledge.
Sufficient information for unanimously accepting only the core payoff vectors Theorem.
Fix a set $N$ of players. Define $\mathbb{G}(N,\cdot)$ to be the set of all TU games $G=(N,v)$ (with $v(S)\in\mathbb{N}_{0}$ for each $S\subseteq N$, as it is always assumed). For each $i\in N$, let $\mathcal{S}_{i}\subseteq\mathcal{N}_{i}$.
  • (i)
    If $\cup_{i\in N}\mathcal{S}_{i}=2^{N}-\{\emptyset\}$, then for each $G\in\mathbb{G}(N,\cdot)$ and each $x\in\mathbb{N}^{N}$, $x\in C(G)$ if and only if $\vdash\Gamma^{G}(\mathcal{S}_{i})\Rightarrow C_{i}(x)$ for all $i\in N$.
  • (ii)
    If $\cup_{i\in N}\mathcal{S}_{i}\subsetneq 2^{N}-\{\emptyset\}$, then there is some $G\in\mathbb{G}(N,\cdot)$ and some $x\in\mathbb{N}^{N}$ such that $x\not\in C(G)$ but $\vdash\Gamma^{G}(\mathcal{S}_{i})\Rightarrow C_{i}(x)$ for all $i\in N$.
Theorem 4.5 resolves Q3 by establishing a non-trivial information requirement for decentralized core computation. Part (i) reveals that unanimous acceptance coincides with core membership precisely when information about coalitions and their values (defined by feasible vectors) are distributed such that every coalition has at least one informed member. Part (ii) demonstrates this condition’s necessity with respect to $\mathbb{G}(N,\cdot)$: any information gap creates vulnerability where non-core vectors may gain unanimous acceptance.
The theorem’s significance extends beyond TU games to any solution concept requiring “coalitional (or even “quantal”) information”—including cores of NTU games (i.e., games without transferable utility). The fundamental insight concerns the relationship between information distribution and achievable normative standard (for example, “fairness” or “justice”, as the core is used to be interpreted): decentralized reasoning faces inherent barriers to implementing fair outcomes, even in moderately-sized games, when information is incomplete.
However, Theorem 4.5.(ii) does not assert that violations of the condition always lead to faulty judgments. Its conclusion is existential: when the coverage condition fails, there exists at least one game where fairness fails to emerge. But not all games are vulnerable. For instance, consider a simple two two-player game $(N,v)$ with $v(\{1\})=0$, $v(\{2\})=v(\{1,2\})=1$. Here, player 1’s knowledge of $v(\{1\})$ proves irrelevant for achieving core selection $(0,1)$, provided player 2 knows $v(\{2\})$ and either player knows $v(\{1,2\})$. [^16] Occasionally, false positives arise due to the discreteness of the model. For example, consider a three-player game where $v(S)=3$ for every $S$ containing agent 1, and $v(S)=0$ otherwise. Suppose agent 1 does not know v(1)v({1}) but knows the value of all other coalitions that include her. In this case, she may accept the payoff vector $(2,1,0)$: because transfers are restricted to integers, no feasible coalition she knows offers her a strictly better allocation. However, this issue is not fundamental. If we allow redistribution in smaller units—e.g., the proposed payoffs are in euros, but agents can transfer cents among themselves—then $(3,0,0)$ becomes the only unanimously acceptable vector, even though agent 1 does not know $v(\{1\})$.
What this suggests is that not all coalitions are equally critical. Some coalitions are functionally relevant—their value must be known to prevent incorrect consensus on unfair allocations—while others are inert. To capture the tightness of our informational requirement, we must identify those coalitions that are functionally necessary for decentralized rejection of non-core outcomes. This motivates the next section, where we apply our proof-theoretic framework to a classic result: the Debreu–Scarf Theorem. There, we will see that the number of functionally necessary coalitions an agent must be informed about grows rapidly with the size of the game—revealing a deeper cognitive barrier to decentralized decision-making.
5 Informational Inconsistency Behind Debreu-Scarf Theorem
The class theorem of Debreu and Scarf (1963) famously states that, as the number of participants in a replica economy grows without bound, the core of the economy converges to the set of competitive equilibria. In this section, we revisit this celebrated result through the lens of our proof-theoretic framework developed in Sections 3 and 4. Our analysis reveals a cognitive inconsistency behind the convergence: as the economy expands, the amount of coalition-related information required for agents to unanimously accept only core allocations grows exponentially; in contrast, competitive equilibria demand a fixed and limited cognitive load—namely, each agent only needs to know prices and her own preferences, regardless of population size. This asymmetry raises foundational questions about the informational plausibility of the core–equilibrium convergence in decentralized settings.
5.1 Pure exchange economy, cooperative game, and $k$-fold replica
A pure exchange economy is a tuple $E=(N,D,\{e_{i}\}_{i\in N},\{u_{i}\}_{i\in N})$, where $N$ is the set of participants of the economy, $D$ is the set of commodities; for each $i\in N$, $e_{i}\in[0,\infty)^{D}$ is her endowment bundle of commodities and $u_{i}:[0,\infty)^{M}\rightarrow\mathbb{R}$ is her utility function. A competitive equilibrium of $E$ is a pair $(p,x)$ of price vector $p\in[0,\infty)^{D}$ and allocation vector $x=(x_{i})_{i\in N}\in\left([0,\infty)^{D}\right)^{N}$ such that $\sum_{i\in N}x_{i}=\sum_{i\in N}e_{i}$, and, for each $i\in N$, $p\cdot x_{i}\leq p\cdot e_{i}$, and for each $y_{i}$ with $p\cdot y_{i}\leq p\cdot e_{i}$, $u_{i}(x_{i})\geq_{i}u_{i}(y_{i})$. A vector $(x_{i})_{i\in N}$ is called an equilibrium allocation of $E$ iff there is some $p\in\mathbb{R}^{D}$ such that $(p,(x_{i})_{i\in N})$ is a competitive equilibrium of $E$.
In the literature, the cooperative game derived from a pure exchange economy is not a TU game but an NTU game. Nevertheless, the framework developed we have developed can also be applied because, as mentioned at the end of Section 4, the core in NTU game concerns only the information about coalitions. Given $E=(N,D,\{e_{i}\}_{i\in N},\{u_{i}\}_{i\in N})$, we derive a NTU game $G(E)=(N,V)$ where for each $S\subseteq N$, $V(S):=\{x=(x_{i})_{i\in N}\in\left([0,\infty)^{D}\right)^{N}:\sum_{i\in S}x_{i}\leq\sum_{i\in S}e_{i}\}$. An allocation vector $x$ is called feasible for coalition $S$ iff $x\in V(S)$.
For coalition $S\subseteq N$ and allocation vectors $x$ and $y$, we say that $y$ dominates $x$ via $S$, denoted by $y$ domS $x,$ iff the following two conditions hold:
  • (i)
    $u_{i}(y_{i})\geq_{i}u_{i}(x_{i})$ for each $i\in S$;
  • (ii)
    $u_{i}(y_{i})>_{i}u_{i}(x_{i})$ for some $i\in S$.
[…]
Without loss of generality, in this section we focus on the simplest 2-person-2-commodity pure exchange economy (i.e., Edgeworth Economy) studies in Shapley (1975) where
  • E1.
    $e_{1}=(1,0)$ and $e_{2}=(0,1)$;
  • E2.
    $u_{1}=u_{2}$;
  • E3.
    $u_{i}$ is continuous, increasing, strictly concave, and homothetic
As shown in Shapley (1975), this economy has a unique equilibrium allocation; the core of the corresponding NTU game can be visualized using the classic Edgeworth box: it is the segment along the contract curve (the diagonal) between the indifference curves of the two agents that pass through the initial endowment. While the equilibrium allocation lies within this core, the core includes many other feasible allocations. Debreu–Scarf Theorem shows that as we replicate the economy by adding more and more identical copies of the original agents, the core of the enlarged (replica) economy shrinks, eventually converging to the singleton consisting solely of the equilibrium allocation.
Formally, given $E=(N,D,\{e_{i}\}_{i\in N},\{u_{i}\}_{i\in N})$ and $k\in\mathbb{N}_{+},$ the $k$-fold replica of $E$ is an economy $E_{k}=(N_{k},D,$ $\{e_{i}\}_{i\in N},\{u_{i}\}_{i\in N})$ where $N_{k}=\cup_{i\in N}\{i_{1},...,i_{k}\}$ and for each $i_{t}$ ($t\in\{1,...,k\}$), $e_{i_{t}}=e_{i}$ and $u_{i_{t}}=u_{i}$ in $E$. Hence, each $i_{t}$ can be regarded as a “replica” of $i$ in $E$, and for $E_{k}$, each $i$ is a type.
Debreu-Scarf Theorem.
$\lim_{k\rightarrow\infty}C(E_{k})=\{x^{E}\}$, where $x^{E}$ is the unique equilibrium allocation of $E$.
5.2 The core of replica economy in the lens of information sufficiency
The Debreu–Scarf Theorem implies that, as the number of participants in an economy grows, the competitive equilibrium allocation emerges as the unique outcome immune to coalition deviations. Intuitively, a larger economy admits more possible coalitions, offering greater opportunities for deviation and thus eliminating more allocations from the $C(E)$. Yet from the perspective of decentralized reasoning, this convergence is not without cost: rejecting allocations out of the reduced core requires agents to assess a rapidly expanding set of cooperation opportunities. This brings into focus a critical question of information sufficiency—namely, what coalition-level information is minimally required to ensure that only core allocations are unanimously accepted. In this subsection, we provide a formal analysis of this trade-off between coalitional richness and the informational demands placed on agents in large economies.
First, as mentioned in Section 2, we have to expand the language. To accommodate a non-trivial discussion and, at the same time, to maintain the countability of the language, we have to modify $\textsf{At}_{\text{fv}}$ and $\textsf{At}_{\text{wp}}$ to include also vectors with rational coordinates. Formally, a symbol in the form of $x^{S}$ is legitimate iff $x\in\left((\mathbb{Q}\cap[0,\infty))^{D}\right)^{S}$.
A key complication arises from our commitment to finitism: since proofs must have finite antecedents, we can no longer use the full set of “feasible vectors” (i.e., symbols in the form of $x^{S}$ with $x$ is feasible for $S$) to represent coalition-level knowledge, as we did in Section 4. Nonetheless, as shown in Section 4, since we will still adopt (a modified version of) $\textsf{C}_{i}$-acceptability as the decision criterion, our analysis requires a way to express information about coalitions that allows us to derive implications about the feasibility of allocations. To address this, we introduce a set of new atomic formulae:
$\textsf{At}_{\text{co}}:=\{\gamma(\mathcal{S}):\mathcal{S}\subseteq N_{k}\text{ for some }k\in\mathbb{N}\}$
and a new logical axioms:
  • $\gamma(\mathcal{S})\Rightarrow x^{S}$: $\gamma(\mathcal{S})\in\textsf{At}_{\text{co}},S\in\mathcal{S},\text{ and }y\text{ is feasible for }S$;
  • $\gamma(\mathcal{S})\Rightarrow\lnot x^{S}$: Otherwise.
Intuitively, each $\gamma(\mathcal{S})$ indicates the information about coalitions in $\mathcal{S}$. It can be seen as a “shorthand symbol” for the set $\Gamma(\mathcal{S})$ in Section 4. The new logical axioms equips agents with ability of computing feasibility: if an agent has information about coalitions in $\mathcal{S}$, she is able to derive which allocation is feasible for each coalition in $\mathcal{S}$.
Finally, we need to modify the criterion expressed by (CORE-CRIT). Since now we allow vectors with rational coordinates, the set of all “reasonable” vectors denoted by $\textsf{PV}(G)$ is now infinite, and, due to our commitment to finitism, (CORE-CRIT) is not a legitimate formula anymore. Instead, we replace $\textsf{PV}(G)$ in (CORE-CRIT) by a finite set AL of allocations with rational coordinates:
$\lnot\left(\bigvee\limits_{S\in\mathcal{N}_{i}}\bigvee\limits_{y\in\mathsf{AL}}(y^{S}\wedge(y\geq_{S}x)\wedge(y>_{\{i\}}x)\right)$ (CORE-FINITE)
We use $C^{\textsf{AL}}_{i}(x)$ to denote the formula in CORE-FINITE. We now say that $x$ is $\textsf{C}_{i}^{*}$-acceptable for some finite set of formulae $\Gamma$ iff $\vdash\Gamma\Rightarrow C^{\textsf{AL}}_{i}(x)$ for each finite set AL of allocations with rational coordinates.
Similar to Proposition 4.4, (unanimously) accepting the unique equilibrium allocation does not need any specific information. The point, as noted in Section 4, is how much information is needed for at least one agent to reject other allocations, and this depends upon which coalitions are “functional”. For the $1$-replica (which is equivalent to $E$), it is easy to see that each non-empty coalition is need. In general, in the $k$-replica $E_{k}$ of $E$ ($k>1$), as shown in Shapley (1975), in addition to the “functional” sets in $E_{k-1}$, the following coalition needs to be known for deleting allocations in $E_{k}$:
  • $\{i_{k}\}$ for $i=1,2$;
  • $\{i_{k},j_{t}\}$ for t=1,,kt=1,...,k and $i\neq j$;
  • All sets $S\in N_{k}$ with $|S\cup N_{k}^{i}|=k$ and $|S\cup N_{k}^{i}|=k-1$, $i\neq j$;
  • $N_{k}$.
Here, for each $i\in N$, $N_{k}^{i}$ is the set of players in $E_{k}$ of type $i$ (i.e., replicas of $i$ in $E$). Each coalition in listed above must be known to at least one of its members in order to refute a non-core allocation. As the number of replicas grows, the total number of such coalitions increases rapidly, and so does the average informational load required to eliminate all non-core allocations. In effect, ensuring unanimous rejection of every unfair distribution demands exponentially more information to be distributed across the population.
This stands in sharp contrast to the competitive framework. There, each agent needs only minimal, fixed information: her own preferences and prevailing prices. Crucially, this cognitive requirement remains constant regardless of the size of the economy.
This reveals a fundamental cognitive asymmetry at the heart of the Debreu–Scarf Theorem. While the theorem assures us that, as the number of participants grows, the core converges to the set of competitive allocations, if the decision is made in a decentralized way, as it is assumed for competitive market, the information structure required to sustain this convergence through individual reasoning becomes increasingly problematic. Competitive equilibrium assumes a stable and limited cognitive burden, whereas the cooperative path to the same outcome demands ever-expanding—and ultimately infeasible—cognitive effort. This tension raises critical questions about the plausibility of interpreting Debreu-Scarf Theorem.
Appendix: Proofs
Cancellation of negation Lemma.
[…]
Proof.
$\frac{\frac{\vdots}{\Gamma\Rightarrow\lnot A}\text{ \ \ \ \ }\frac{A\Rightarrow A}{\lnot A,A\rightarrow\text{ \ \ \ }}(\lnot\text{-Left})}{\Gamma,A\Rightarrow\text{ \ \ \ }}(\text{Cut})$
Proof of Proposition 4.3 Proof.
Fix $x\in\mathbb{N}^{N}$. First, since non-logical axioms are well defined (i.e., consistent) because they are based on numerical comparison on natural numbers, it is clear that for each $S\in\mathcal{N}_{i}$ and eachg $y\in\mathsf{PV}(G)$, either  $\vdash\text{ \ \ }\Rightarrow(y\geq_{S}x)\wedge(y>_{\{i\}}x)$ or  $\vdash\text{ \ \ }\Rightarrow\lnot\left((y\geq_{S}x)\wedge(y>_{\{i\}}x)\right)$. We discuss two cases.
[…]
$\frac{\frac{\frac{\vdots}{(y\geq_{S}x)\wedge(y>_{\{i\}}x)\Rightarrow\text{ \ \ }}}{y^{S}\wedge(y\geq_{S}x)\wedge(y>_{\{i\}}x)\Rightarrow\text{ \ \ }}(\wedge\text{-Left})}{\Gamma_{i},y^{S}\wedge(y\geq_{S}x)\wedge(y>_{\{i\}}x)\Rightarrow\text{ \ \ }}(\text{Weakening})$
[…]
Case 2. For some $S\in\mathcal{N}_{1}$ and some $y\in\textsf{PV}(G)$, $\vdash(y\geq_{S}x)\wedge(y>_{\{i\}}x)$. We define $D(x):=\{y^{S}\in\textsf{At}_{\text{fv}}:S\in\mathcal{N}_{i}$ and  $\vdash\text{ \ \ }\Rightarrow(y\geq_{S}x)\wedge(y>_{\{i\}}x)\}$. Here we consider two subcases.
[…]
$\frac{\frac{\vdots}{\Gamma_{i},y^{S}\Rightarrow\text{ \ \ }}}{\Gamma_{i},y^{S}\wedge(y\geq_{S}x)\wedge(y>_{\{i\}}x)\Rightarrow\text{ \ \ }}(\wedge\text{-Left})$
[…]
$\frac{\frac{\frac{y^{S}\Rightarrow y^{S}}{\Gamma_{i}\Rightarrow y^{S}}(\text{Weakening})\text{ \ \ }\frac{\vdots}{\Gamma_{i}\Rightarrow(y\geq_{S}x)\wedge(y>_{\{i\}}x)}}{\Gamma_{i}\Rightarrow(y^{S}\wedge(y\geq_{S}x)\wedge(y>_{\{i\}}x)}(\wedge\text{-Right})}{\Gamma_{i}\Rightarrow\bigvee\limits_{S\in\mathcal{N}_{i}}\bigvee\limits_{y\in\mathsf{PV}(G)}(y^{S}\wedge(y\geq_{S}x)\wedge(y>_{\{i\}}x)}(\vee\text{-Right})$
That is, $\vdash\Gamma_{i}\Rightarrow\lnot C_{i}(x)$.
That not both $\vdash\Gamma_{i}\rightarrow C_{i}(x)$ and $\vdash\Gamma_{i}\rightarrow\lnot C_{i}(x)$ hold simultaneously can be seen clearly from the proof, because all cases need “for all” statement (or, as in Case 2.2, its negation “for some”), which cannot be both true and false.
Proof of Proposition 4.4 Proof.
Suppose that $\nvdash\Gamma^{G}(\mathcal{S})\Rightarrow\lnot C_{i}(x)$ for some $S\in 2^{N}$. Since $\Gamma^{G}(\mathcal{S})$ satisfies condition (D) in Proposition 4.3, it follows that $\vdash\Gamma^{G}(\mathbb{\mathcal{S}})\Rightarrow\lnot C_{i}(x)$. By reflecting on the inference rules, this implies that for some $S\in\mathcal{n}_{i}$ and $y\in\textsf{At}_{\text{fv}}$, $\vdash\Gamma^{G}(\mathcal{S})\Rightarrow y^{S}\wedge y\geq_{S}x\wedge y>_{\{i\}}x$, which follows that $\vdash\Gamma^{G}(\mathcal{S})\Rightarrow y^{S}$, $\vdash\Gamma^{G}(\mathcal{S})\Rightarrow y\geq_{S}x$, and $\vdash\Gamma^{G}(\mathcal{S})\Rightarrow y>_{\{i\}}x$. By $\vdash\Gamma^{G}(\mathcal{S})\Rightarrow y^{S}$, it follows that $\sum_{j\in S}y_{j}\leq S$. Then, by the definition of the core (Definition 1, Coalitional Stability), (1) $x_{i}\geq y_{i}$, which, by our definition of the non-logical axioms, implies  $\vdash\text{ \ \ }\Rightarrow\lnot y>_{\{i\}}x)$, and, consequently, we obtain $\vdash\Gamma^{G}(\mathcal{S})\Rightarrow\lnot(y\geq_{S}x)$ by Weakening. Yet this is in conflict with $\vdash\Gamma^{G}(\mathcal{S})\Rightarrow y\geq_{S}x$, an implication to our (contradictory) assumption. Since our system is consistent, this is impossible. Therefore, $\nvdash\Gamma^{G}_{i}(\mathcal{S})\Rightarrow C_{i}(x)$ for all $\mathcal{S}\subseteq 2^{N}$.
Proof of Proposition 1 Proof.
The Only-if part holds by Weakening rule. For the If part, we show by contrapositive. Suppose that $\nvdash\Gamma^{G}(\mathcal{S}\cup\mathcal{T})\Rightarrow C(x)$. Since $\Gamma^{G}(\mathcal{S}\cup\mathcal{T})$ satisfies conditions D1 and D2 in Proposition 4.3, it follows that $\vdash\Gamma^{G}(\mathcal{S}\cup\mathcal{T})\Rightarrow\lnot C(x)$, which means that for some $y\in\textsf{PV}(G)$ and $S\in\mathcal{N}_{i}$, $\vdash\Gamma^{G}(\mathcal{S}\cup\mathcal{T})\Rightarrow y^{S}\wedge(y\geq_{S}x)\wedge(y>_{\{i\}}x)$. To imply both $(y\geq_{S}x)$ and $(y>_{\{i\}}x)$ only needs non-logical axioms, which does not matter which formulae are in the antecedent. Hence the critical implication is that $\vdash\Gamma^{G}(\mathcal{S}\cup\mathcal{T})\Rightarrow y^{S}$. Since $S\in\mathcal{N}_{i}$ and $\mathcal{T}\cap\mathcal{N}_{i}=\emptyset$, it follows that $S\in\mathcal{S}$, which follows that $y^{S}\in\Gamma^{G}(\mathcal{S})$. Therefore, $\vdash\Gamma^{G}(\mathcal{S})\Rightarrow\lnot C(x)$, that is, $\nvdash\Gamma^{G}(\mathcal{S})\Rightarrow C(x)$.
Proof of Theorem 4.5 Proof.
(i) (If) We show this statement by contrapositive. Suppose that $x\not\in C(G)$. Then there is some non-empty $S\subseteq N$ such that $\sum_{j\in S}x_{j}<v(S)$, which, since $x\in\mathbb{N}^{N}$ and $v(S)\in\mathbb{N}$, follows that $\sum_{j\in S}x_{j}<v(S)-1$. Since ${S}_{j}\subseteq\mathcal{N}_{i}$ for each $j\in N$ and $\cup_{j\in N}\mathcal{S}_{j}=2^{N}-\{\emptyset\}$, there is some $i\in N$ such that $i\in S$ and $S\in\mathcal{S}_{i}$. Consider a vector $y\in\mathbb{N}^{N}$ defined as follows:
$y_{j}=\begin{cases}x_{j}+1&\text{if }j=i,\\ x_{j}&\text{otherwise}.\end{cases}$
By definition, it is clear that $y^{S}\in\Gamma^{G}(\mathcal{S}_{i})$, and consequently it follows that $\vdash\Gamma^{G}(\mathcal{S}_{i})\Rightarrow\lnot C_{i}(x)$. Since $\Gamma^{G}(\mathcal{S}_{i})$ satisfies condition (D) in Proposition 4.3, it follows that $\nvdash\Gamma^{G}(\mathcal{S}_{i})\Rightarrow C_{i}(x)$.
[…]
(ii) Let $S$ be a non-empty subset of $N$ such that $S\notin\mathcal{S}_{i}$ for each $i\in S$. We define $G=(N,v)\in\mathcal{G}(N,\cdot)$ such that $v(N)=v(S)=|N|$ and $v(T)=0$ for all other $T\subseteq N$. To show the statement, we consider the following two cases.
[…]
Case 2. $S\subsetneq N$. It can be seen that $C(G)\cap\mathbb{N}^{N}=\{z\in\mathbb{N}^{N}:\sum_{j\in S}z_{j}=|N|$ and $z_{j}=0$ for each $j\notin S\}$. Consider x=(1,,1).x=(1,...,1). Though $x\notin C(G)$, in a manner similar to that in Case 1, we show that $\vdash\Gamma^{G}(\mathcal{S}_{i})\Rightarrow C_{i}(x)$ for each $i\in N$, no matter whether $i$ knows $v(N)$ (i.e, $N\in\mathcal{S}_{i}$) or not.